THE STRUCTURE EFFECT OF AU AND PT CLUSTERS ON THE VIBRATIONAL SPECTRA OF AD-SORBED METHANE MOLECULES

\underline{LI XIAO}, LICHANG WANG, Department of Chemistry and Biochemistry, Southern Illinois University, Carbondale, IL 62901; ,.

The adsorption of CH_{4} molecules on different Au and Pt clusters was investigated using density functional theory calculations. The results show that the vibrational spectra of CH_{4} depend strongly on the adsorption site and orientation of CH_{4} as well as the structure of Au and Pt isomers. Our results also indicate a possible increase of the peak temperature with cluster size and the broadening of the spectra that could be observed in the temperature-programmed desorption experiment.